OpenStudy (anonymous):

Find (x^2+1)/x if ...

OpenStudy (anonymous):

\[\sqrt[3]{x^{2}}-3\sqrt[3]{x}+1=0\]

OpenStudy (mathmath333):

try to put \(x^{1/3}=a\) and solve quadratic

OpenStudy (anonymous):

i tried that but i could not get any result ??

OpenStudy (mathmath333):

that method looks cumbersome anyway

OpenStudy (triciaal):